# DRP: Distilled Reasoning Pruning with Skill-aware Step Decomposition for Efficient Large Reasoning Models

### v2

@@ 6.3 Influence of Teacher Model Choice
To evaluate the sensitivity of our pruning framework to the choice of teacher model, we experiment with three different large language models (LLMs): GPT-4o, Gemini 2.0 Flash, and ChatGPT. As shown in Table 3, all teacher models lead to consistent improvements in both accuracy and token efficiency compared to the base model, confirming the robustness and generalizability of our approach.
Among them, GPT-4o yields the strongest performance across most benchmarks, particularly in token compression. However, the overall differences between teacher models remain modest, suggesting that while stronger teachers can produce slightly more optimal pruning decisions, our method remains effective even when relying on weaker or more accessible models.
This indicates that our framework does not overfit to a specific teacher’s internal reasoning patterns, and can generalize pruning supervision across a variety of LLMs. Notably, we observe that variations across teachers are more pronounced in token usage than in accuracy—highlighting that our method’s core benefit, skill-level reasoning distillation, remains stable regardless of teacher strength.
@@ Appendix D Skill Base Decomposition Evaluation
To assess the quality of our skill-based step decomposition, we conduct a pairwise comparison experiment using Gemini 2.0 Flash as the judge Gao et al. (2023); Li et al. (2024, 2025a). For each reasoning trace $T$, we generate two segmentations: (1) our proposed skill-based decomposition, and (2) a baseline sentence-based split (obtained by naïvely splitting at punctuation or sentence boundaries).

### v3

@@ 6.3 Influence of Teacher Model Choice
To evaluate the sensitivity of our pruning framework to the choice of teacher model, we experiment with four different large language models (LLMs): GPT-4o, Gemini 2.0 Flash, ChatGPT, and DeepSeek-V3. As shown in Table 3, all teacher models yield consistent improvements in both accuracy and token efficiency over the base model, confirming the robustness and generalizability of our approach.
Among them, GPT-4o achieves the strongest performance, especially in token compression. Using DeepSeek-V3 as the teacher also leads to clear gains over the student baseline, with results comparable to other teacher models. This suggests that teacher–student family alignment is not critical for the effectiveness of our DRP method.
Overall, while stronger teachers can yield slightly more optimal pruning decisions, the differences remain modest. Our method generalizes well across teacher models and does not overfit to a specific model’s reasoning style. Notably, we observe greater variation in token usage than in accuracy, indicating that our skill-aware pruning retains its core benefits regardless of teacher strength.
@@ Appendix D Skill Base Decomposition Evaluation
To assess the quality of our skill-based step decomposition, we conduct a pairwise comparison experiment using Gemini 2.0 Flash as the judge Gao et al. (2023); Li et al. (2024a, 2025a). For each reasoning trace $T$, we generate two segmentations: (1) our proposed skill-based decomposition, and (2) a baseline sentence-based split (obtained by naïvely splitting at punctuation or sentence boundaries).